If the binding energy per nucleon in 3Li7 and 2He4 nuclei are respectively 5.60 MeV and 7.06 Me V, then the energy of proton in the reaction 3Li7 + p --->2 2He4 is


        (a)    19.6 MeV          


(b)    2.4 MeV


        (c)    8.4 MeV


        (d)    17.3 MeV


Grade:upto college level

3 Answers

Amit Saxena
10 years ago

(d)

Applying principle of energy conservation, Energy of proton

=      total B.E. of 2a - energy ofLi7

=      8 × 7.06 - 7 × 5.6

=      56.48 - 39.2 = 17.28 MeV

Anurag Thamke
5 years ago
By applying principle ofel energy conservation,   
Total binding energy of product -energy of lithium 
=8×7.60 - 7×5.6
=56.48-) - 39.2 =17.28
 
 
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Q=2(4×7.06)−7×(5.60)
=(56.48−39.2)MeV
=17.28MeV
=17.3MeV

Thanks and Regards